# NoveltyRank: Estimating Conceptual Novelty of AI Papers

### v1

@@
NoveltyRank:
Estimating Conceptual Novelty of AI Papers
Zhengxu Yan1,*, Han Li2,*, and Yuming Feng2,*
1Department of Computer Science, Stanford University
2Department of Electrical Engineering, Stanford University
jasonyan@stanford.edu, cli27@stanford.edu, yumingf@stanford.edu
@@ 1 Introduction
With the growing ease of academic publishing, the volume of research papers, especially in AI-related fields, has surged dramatically. This "flood" of publications make it difficult for truly novel and impactful work to stand out, and manual novelty assessment is often unstable and time-consuming. Our project aims to develop a model that estimates and ranks the conceptual novelty of AI papers, enabling a data-driven and scalable assessment of research originality. Such a system can help researchers efficiently identifying submissions that introduce genuinely innovative ideas rather than minor variants, and provide conference reviewers with a quantitative and consistent signal of novelty.
Our approach evaluates novelty primarily through a paper’s title, abstract, and semantic similarity to prior literature. Given the motivation of novelty estimation, we explore two task formulations with different modeling objectives, each offering a different perspective: (1) Binary classifcation, which predicts the paper’s absolute novelty from learned patterns of prior novel works, and (2) Pairwise novelty comparison, which learns to distinguish papers by relative novelty over others. We fine-tune Qwen3-4B-Instruct-2507 qwen and SciBERTscibert on both tasks, benchmarking against GPT-5.1 to analyze how task formulation and modeling choices affect performance. The implementation is publicly available at https://github.com/ZhengxuYan/NoveltyRank.
2 Dataset and Features
2.1 Description
@@
Our dataset is built from web-scraped data and the public ICLR 2017–2025 dataset gonzalez2024learning, totaling 60,294 papers published between 2023 and 2025: 50,442 randomly sampled arXiv papers and 9,852 accepted by top-tier venues across six domains (AI, ML, CV, Robotics, NLP, and Cryptography). Each entry includes metadata such as paper ID, publication date, authors, title, and abstract.
Conference acceptance serves as a proxy for novelty, assuming accepted papers generally reflect more original or impactful contributions: Accepted papers are labeled as positives (1), while randomly sampled arXiv papers are negatives (0). To prevent temporal leakage, we split data by publication date: Training on papers from 2024 to early 2025, and testing on papers after March 15, 2025. This setup ensures evaluation on future, unseen publications, aligning with real-world scenarios.
2.2 Embedding
We embed each paper’s title and abstract using SPECTER2 specter2, a transformer fine-tuned for scientific document representation. We employed two adapter modules:
  1. 1.
    Classification Embedding – captures a paper’s semantic identity, such as topic, technical depth, and conceptual contribution.
  2. 2.
    Proximity Embedding – trained with citation-based contrastive learning to measure closeness in citation space. Based on proximity embedding, we adopt Faiss faiss for scalable nearest-neighbor search and compute papers’ similarity scores given their top-10 similar prior literature (rolling by publication date), serving as additional engineered features.
Combining these embeddings enables the model to assess both content-level representation and contextual similarity for novelty evaluation.
@@ 3 Models
To implement the NoveltyRank framework, we formulate novelty estimation through two distinct task both aligned with the goal of evaluating scientific innovation: Binary Classification (evaluating the absolute novelty of a single paper) and Pairwise Comparison (assessing relative novelty between two papers). These tasks differ in input structure and supervision.
To benchmark performance across computational scales, we experiment each task formulation on three models listed below following a specific logic: as model size decreases, the degree of task-specific adaptation increases. Such comparison determine whether smaller models with targeted parameter updates can match or surpass larger general-purpose models, verifying the feasibility of lightweight deployment.
@@ 4 Task Formulation 1: Binary Novelty Classification
For binary novelty classification, we formulate the task as a supervised learning problem where the model primarily use papers’ title, abstract, and similarity scores to predict binary novelty label (0 or 1). The objective is to learn general patterns and absolute criteria from established novel papers to assess the novelty of unseen samples.
@@ 4.1 Qwen3-4B: SFT and DPO Fine-tuning
The full pipeline of Qwen3-4B model for binary novelty classification task is illustrated in Fig. 1. Because novelty is inherently comparative, we augment this input with a Similarity Report generated by an LLM Judge, which summarizes overlaps and distinctive contributions relative to the top-$K$ similar papers retrieved via FAISS faiss. This comparison serves provides an additional signal for decision-making process.
@@
Supervised Fine-Tuning (SFT)
In SFT, we apply cross-entropy loss between the model’s generated tokens and ground-truth labels. This setup naturally yields a binary supervision signal, training the model to generate discrete outputs (0 or 1). We avoid prompting for continuous confidence scores (e.g., 0.85), as LLMs generate text tokens rather than mathematically grounded probabilities. The decimals ouputs would be linguistic hallucinations, rendering the model unreliable for quantitative confidence estimation.
Direct Preference Optimization (DPO)
@@ 4.2 Fine-tuned SciBert
These features are fused into a multi-modal representation and passed through a custom classification head consisting of a three-layer feed-forward network ($2306\to 512\to 128\to 2$) with ReLU activation and a dropout rate of 0.1. The model is trained using Cross-Entropy loss with an AdamW optimizer ($\eta=2e^{-5}$) and a linear warmup scheduler for 5 epochs.
4.3 Evaluation Metrics
@@ 4.4 Results and Discussion
Table 1: Test Performance of Binary Classification
@@
Despite higher accuracy (e.g., SciBERT’s 74.4%), fine-tuned models suffer from dataset imbalance. Low recall (especially SciBERT’s 0.313) indicates they models collapse into conservative classifiers, minimizing loss by over-predicting the majority "non-novel" class.
@@ 5.1 Qwen3-4B: SFT and DPO Fine-tuning
The same prompt-engineering methods and the SFT and DPO methods described in Section 4.1 applies here. However, the supervision signal shifts from binary labels (0 or 1) to positional indicators (Paper A or Paper B). This allows the model to adapt the same optimization pipeline to a relative comparison setting.
@@ 5.2 Siamese SciBERT Network
5.3 Evaluation Metrics
@@ 5.4 Results and Discussion
Table 2: Performance of Pairwise Comparison
@@ 6 Conclusion
This project introduces NoveltyRank, a system to evaluatw the conceptual originality of AI papers. We find that pairwise comparison formulation offers a significantly cleaner and more effective learning signal than binary classification formulation. Moreover, We investigate three model scales and find that the cost-effective, domain-specific SciBERT, fine-tuned under the comparison setting, achieves best results. This demonstrates that comparative structure with small-model fine-tuning provides a more efficient and effective solution than increasing model size.
@@ Acknowledgments
We would like to express our sincere gratitude to Cody Ho and Ryan Lian for their helpful feedback and guidance during the course of this project.
@@
A. Deployment
The best-performed model is deployed in a user-friendly web application [^1] https://novelty-rank.vercel.app/ that allows interactive exploration of novelty evaluation and ranking.
@@ B. Hyperparameters
Key hyperparameters for Qwen3-4B (SFT and DPO) and SciBERT includes:

### v2

# NoveltyRank: A Retrieval-Augmented Framework for Conceptual Novelty Estimation in AI Research

@@
Abstract
The accelerating pace of scientific publication makes it difficult to identify truly original research among incremental work. We propose a framework for estimating the conceptual novelty of research papers by combining semantic representation learning with retrieval-based comparison against prior literature. We model novelty as both a binary classification task (novel vs. non-novel) and a pairwise ranking task (comparative novelty), enabling absolute and relative assessments. Experiments benchmark three model scales, ranging from compact domain-specific encoders to a zero-shot frontier model. Results show that fine-tuned lightweight models outperform larger zero-shot models despite their smaller parameter count, indicating that task-specific supervision matters more than scale for conceptual novelty estimation. We further deploy the best-performing model as an online system for public interaction and real-time novelty scoring.
@@ 1 Introduction
The volume of research publications, particularly in AI-related fields, has accelerated dramatically due to the accessibility of modern academic workflows. This surge has made it increasingly difficult for genuinely novel work to stand out, as incremental papers often blend into the growing literature. Manual novelty assessment is time-consuming, subjective, and difficult to scale, motivating automated methods for estimating the originality of research ideas. Our goal is to develop a model that estimates and ranks the conceptual novelty of AI research papers, providing a data-driven, consistent signal of originality. Such a system may help identify unconventional research directions and highlight submissions that introduce genuinely new ideas rather than minor variations.
We evaluate conceptual novelty using semantic information from a paper’s title and abstract, along with its similarity to prior literature. To operationalize this, we explore two task formulations: (1) binary classification, which predicts absolute novelty from supervised examples, and (2) pairwise comparison, which learns relative novelty through pairwise preference signals. We fine-tune Qwen3-4B-Instruct-2507 qwen and SciBERT scibert on both tasks, and benchmark against GPT-5.1 openai_gpt51 in a zero-shot setting to analyze the impact of model scale and supervision.
Our contributions are threefold: (1) we formalize conceptual novelty estimation as context-aware conceptual deviation from prior literature and instantiate it through binary classification and pairwise comparison tasks; (2) we benchmark domain-specific, mid-sized fine-tuned, and zero-shot frontier models, finding that targeted fine-tuning of compact models outperforms zero-shot usage of significantly larger models; (3) we deploy the best-performing model as an interactive system for real-time novelty scoring and retrieval, demonstrating practical usability for literature exploration.
Code is available on GitHub [^1] https://github.com/ZhengxuYan/NoveltyRank, and the system is deployed as a web application [^2] https://novelty-rank.vercel.app/ for interactive exploration and community feedback.
2 Related Work
Document Representation and Scholarly Retrieval. Transformer-based representations have become central to scientific document indexing and retrieval. SPECTER and SPECTER2 map papers into embedding spaces aligned with citation intent, enabling semantic search and clustering specter2. Systems such as arXiv Sanity Preserver karpathy2018arxiv, Semantic Scholar, and topic-based recommenders surface relevant works via semantic similarity or metadata signals. However, these systems emphasize relevance rather than conceptual originality. Our framework is complementary: novelty scores can be layered on top of retrieval pipelines to surface unconventional or frontier ideas within a topic.
Novelty and Originality Estimation. Prior work models novelty as deviation from historical literature via supervised classification ghosal2018tapdlnd10corpus, semantic redundancy detection ghosal-etal-2018-novelty, or network-based atypicality in citation/co-author graphs AMPLAYO2018542. Outlier-centric approaches estimate novelty by density or distance metrics in embedding space, e.g., fastText + LOF for biomedical titles JEON2023101450. These methods rely heavily on citation structures or handcrafted statistical assumptions. In contrast, our approach incorporates transformer-based representations with retrieval-anchored contextual signals, enabling the model to assess context-aware conceptual deviation rather than mere semantic proximity. We further study how task formulation and model scale shape novelty prediction performance.
3 Dataset
3.1 Data Source and Labeling
@@
Our dataset combines web-scraped arXiv entries with the public ICLR 2017–2025 dataset gonzalez2024learning, totaling 60,294 papers published between 2023 and 2025. The corpus includes 50,442 randomly sampled arXiv papers and 9,852 papers accepted to top-tier venues across six domains (AI, ML, CV, Robotics, NLP, and Cryptography). For each entry, we retain metadata including paper ID, publication date, title, authors, and abstract.
Following prior work using venue acceptance as a heuristic signal for originality, we adopt conference acceptance as a proxy label for conceptual novelty: accepted papers are assigned label $1$ (positive) and randomly sampled arXiv papers are assigned label 0 (negative). To prevent temporal leakage, we perform a chronological split: models are trained on papers from 2024 to early 2025 and evaluated on papers published after March 15, 2025. This setup reflects a real-world deployment scenario where novelty estimation is applied to future, unseen submissions.
3.2 Document Representations
We encode each paper using SPECTER2 specter2, a transformer model trained for scientific document representation. Titles and abstracts are mapped to embedding vectors via two model heads with different semantic emphases:
  • Classification Embedding — captures a paper’s semantic content for downstream prediction.
  • Proximity Embedding — optimized with citation-based contrastive learning to reflect relational distance between papers in citation space.
These representations provide complementary signals: the former models internal semantics, while the latter situates the paper within the scientific landscape.
3.3 Neighborhood-based Features (Retrieval-Augmented)
To incorporate contextual signals beyond intrinsic document semantics, we adopt a retrieval-augmented design. Using proximity embeddings, each paper retrieves its top-$10$ most similar prior works via Faiss faiss, restricted to strictly earlier publication dates to avoid future information leakage. From these retrieved neighborhoods, we compute statistical features such as similarity aggregates (e.g., max/mean similarity) and deviation profiles that summarize how atypical a paper is relative to its closest historical neighbors. These retrieval-derived features are concatenated with the base document embeddings, enabling the model to assess novelty through both semantic representation and contextual deviation from prior literature.
@@ 4 Models
To implement the NoveltyRank framework, we formulate novelty estimation through two distinct tasks both aligned with the goal of evaluating scientific innovation: Binary Classification (evaluating the absolute novelty of a single paper) and Pairwise Comparison (assessing relative novelty between two papers). These tasks differ in input structure and supervision.
To benchmark performance across computational scales, we experiment with each task formulation on three models listed below following a specific logic: as model size decreases, the degree of task-specific adaptation increases. Such comparison determines whether smaller models with targeted parameter updates can match or surpass larger general-purpose models, verifying the feasibility of lightweight deployment.
@@
Our models are trained in PyTorch paszke2019pytorch with the HuggingFace Transformers library wolf2019huggingface.
@@ 5 Task Formulation 1: Binary Novelty Classification
For binary novelty classification, we formulate the task as a supervised learning problem where the model primarily uses papers’ title, abstract, and similarity scores to predict binary novelty label (0 or 1). The objective is to learn general patterns and absolute criteria from established novel papers to assess the novelty of unseen samples.
@@ 5.1 Qwen3-4B: SFT and DPO Fine-tuning
The full pipeline of Qwen3-4B model for binary novelty classification task is illustrated in Fig. 1. Because novelty is inherently comparative, we augment this input with a Similarity Report generated by an LLM Judge, which summarizes overlaps and distinctive contributions relative to the top-$K$ similar papers retrieved via FAISS faiss. This comparison provides an additional signal for decision-making process.
@@
SFT
In SFT, we apply cross-entropy loss between the model’s generated tokens and ground-truth labels. This setup naturally yields a binary supervision signal, training the model to generate discrete outputs (0 or 1). We avoid prompting for continuous confidence scores (e.g., 0.85), as LLMs generate text tokens rather than mathematically grounded probabilities. The decimal outputs would be linguistic hallucinations, rendering the model unreliable for quantitative confidence estimation.
DPO
@@
Full hyperparameter configurations for SFT and DPO are provided in Table 3 in Appendix.
@@ 5.2 Fine-tuned SciBERT
These features are fused into a multi-modal representation and passed through a custom classification head consisting of a three-layer feed-forward network ($2306\to 512\to 128\to 2$) with ReLU activation and a dropout rate of 0.1. The model is trained using Cross-Entropy loss with an AdamW optimizer ($\eta=2e^{-5}$) and a linear warmup scheduler for 5 epochs. Full hyperparameter configurations for the SciBERT encoder are provided in Table 4 in Appendix.
5.3 Evaluation Metrics for Binary Novelty Classification
@@ 5.4 Results and Discussion
Table 1: Test Performance of Binary Classification (n=10,889)
@@
Despite higher accuracy (e.g., SciBERT’s 74.4%), fine-tuned models suffer from dataset imbalance. Low recall (especially SciBERT’s 0.313) indicates the models collapse into conservative classifiers, minimizing loss by over-predicting the majority "non-novel" class.
@@ 6.1 Qwen3-4B: SFT and DPO Fine-tuning
The same prompt-engineering methods and the SFT and DPO methods described in Section 5.1 apply here. However, the supervision signal shifts from binary labels (0 or 1) to positional indicators (Paper A or Paper B). This allows the model to adapt the same optimization pipeline to a relative comparison setting.
@@ 6.2 Siamese SciBERT Network
6.3 Evaluation Metrics for Pairwise Novelty Comparison
@@ 6.4 Results and Discussion
Table 2: Performance of Pairwise Comparison (n=9,531)
@@ 7 Conclusion
This project introduces NoveltyRank, a system to evaluate the conceptual originality of AI papers. We find that pairwise comparison formulation offers a significantly cleaner and more effective learning signal than binary classification formulation. Moreover, we investigate three model scales and find that the cost-effective, domain-specific SciBERT, fine-tuned under the comparison setting, achieves best results. This demonstrates that comparative structure with small-model fine-tuning provides a more efficient and effective solution than increasing model size.
@@ Acknowledgments
We would like to express our sincere gratitude to Cody Ho and Ryan Lian for their valuable feedback and guidance throughout this project. We are also grateful to Andrew Ng and Kian Katanforoosh for the insightful course content and discussions that shaped our understanding of applied machine learning. Finally, we thank Thinking Machines Lab for providing access to Tinker compute resources that supported our experiments.
@@
A. Dataset
@@ B. Hyperparameters
Key hyperparameters for Qwen3-4B (SFT and DPO) and SciBERT include: